undeadknight26 (undeadknight26):

The table shows data from a survey about the amount of time high school students spent reading and the amount of time spent watching videos each week (without reading). Reading Video 4 4 4 5 5 6 5 8 5 9 6 10 7 11 8 12 8 14 9 25 Which response best describes outliers in these data sets? Neither data set has suspected outliers. The range of data is too small to identify outliers. Video has a suspected outlier in the 25 hour value. The 25 hour value for video does not pass the outlier test of 1.5 • (IQR) + Q3.

undeadknight26 (undeadknight26):

@hartnn @annas @Agent47 @SolomonZelman @Darry&Gator @hartnn please help.

undeadknight26 (undeadknight26):

Forgot a choice: The 25 hour value for video does not pass the outlier test of 1.5 • (IQR) + Q3.

undeadknight26 (undeadknight26):

I think it is either c or d.

OpenStudy (anonymous):

@agent0smith @ranga

OpenStudy (anonymous):

@mathmale

OpenStudy (ranga):

The reading data does not have any numbers that looks way off. But the video data has 25 that looks way off. To test if 25 is an outlier, first find the first and third quartiles: 4, 5, 6, 8, 9, 10, 11, 12, 14, 25 ^ ^ ^ Q1 M Q3 First Quartile Q1 = 6 Median M = (9+10)/2 = 9.5 Third Quartile Q3 = 12 Inter-Quartile Range (IQR) = Q3 - Q1 = 12 - 6 = 6 Upper Bound for Outliers = Q3 + 1.5 * IQR = 12 + 1.5 * 6 = 12 + 9 = 21. Since 25 lies beyond the upper bound, 25 is an outlier. So the answer is C.